OpenStudy (anonymous):

Determine if the following Serie converges or not \[\Large{\sum_{k\in\mathbb{N}}\frac{(-1)^{k}}{k^{2}+1}}\]

OpenStudy (anonymous):

Alternating series test.

OpenStudy (anonymous):

i know but i couldnt apply it.. can u show me please ? i have exam on wendensday =(

OpenStudy (anonymous):

Well, show the absolute value is always decreasing.

OpenStudy (anonymous):

i think the test is for showing if the serie monoton decreased.. bu

OpenStudy (anonymous):

hmm how to apply it?

OpenStudy (anonymous):

Show \(a_n > a_{n+1}\).

OpenStudy (anonymous):

ok i will make a step but i would be happy if you help me by the next steps..

OpenStudy (anonymous):

Well, the absolute value is always decreasing.

OpenStudy (anonymous):

\[ n<n+1\\ n^2<(n+1)^2\\ n^2+1<(n+1)^2+1\\ \frac{1}{n^2+1}>\frac{1}{(n+1)^2+1}\\ |a_n|>|a_{n+1}| \]

OpenStudy (anonymous):

\[\Large{\sum_{k\in\mathbb{N}}\frac{(-1)^{k}}{k^{2}+1}=(-1)^{k}(\frac{1}{k^{2}+1})}\] and i think \[\Large{ \frac{1}{k^{2}+1}\geq \frac{1}{(k+1)^{2}+1}}\]

OpenStudy (anonymous):

aah ok.. adn how to plugin it to question so that in exam i can get full points for this question?

OpenStudy (anonymous):

Just invoke the alternating series test.

OpenStudy (anonymous):

ok i will try..

OpenStudy (anonymous):

thank you very much